solve using the elimination method 6x-2y=14 and 2x+5y-18
Answer by mananth(16946)

6 x -2 y = 14 .............1
Total value
2 x + 5 y = -18 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 2
30 x -10 y = 70
4 x + 10 y = -36
Add the two equations
34 x = 34
/ 34
x = 1
plug value of x in (1)
6 x -2 y = 14
6 -2 y = 14
-2 y = 14 -6
-2 y = 8
y = -4
x= 1
y= -4
m.ananth@hotmail.ca